# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 5.1 Applying DNN
In practice, we approximate $g_{o}$ by a fully connected ReLU network and model the linear term $\mbox{${\theta}$}^{\top}{\bf Z}$ using a parallel linear layer. The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3). This architecture jointly learns the nuisance function $g$ and a finite–dimensional linear effect of $\bf Z$} in a way that mirrors the model structure.
To evaluate the approximated log-likelihood (3) in practice, we discretize the time axis on a fixed grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$. The dataset is then expanded in the usual counting‐process manner. Each subject $i$ with data $(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})$ is represented by one dummy-observation for every interval $t_{k},t_{k+1})$ such that $t_{k}\leq T_{i}$. All dummy-observations inherit $(\mathbf{X}_{i},\mathbf{Z}_{i})$; the event indicator $\Delta_{ik}=\mathbf{1}\{T_{i}\in(t_{k-1},t_{k}],\Delta=1\}$ is set to 0 except for the final interval containing $T_{i}$, where it takes the original value $\Delta_{i}$. This choice of grid and data expansion enabling efficient computation within standard deep-learning frameworks. The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (therneau2000cox; hu2023conditional).
@@ 5.2 Estimation of the Covariance Matrix $I(\mbox{${\theta}$}_{o})^{-1}$
Inference on $\mbox{${\theta}$}_{0}$ relies an estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$ from Theorem 4, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$. Recall that the efficient information is
@@
where $\bf g$$^{\ast}$ is the event–weighted projection of $\bf Z$} onto the nuisance tangent space. In practice we estimate $\bf g$$^{\ast}$ using an additional set of neural networks and then construct a plug–in estimate of $I(\mbox{${\theta}$}_{o})$ from the corresponding residuals.
Specifically, we approximate $\bf g$$^{\ast}$ coordinatewise. For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by
$\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}\,.$
Each network $j$, $j=1,\ldots,p$, takes $(T_{i},\mathbf{X}_{i})$ as input, $Z_{ij}$ as target, and uses the same architecture and hyperparameters as the main model; empirically, these choices produce stable estimates and are not sensitive to tuning.
To avoid overfitting, we employ 55–fold cross–fitting. The data are split into five folds, for each fold and each coordinate $j$, the corresponding network is trained on the remaining four folds (restricting to $\Delta=1$) and predictions $\widehat{g}^{\ast}_{j}(T_{i},X_{i})$ are obtained for the held-out fold. for all $i$ in the held–out fold. Aggregating across folds yields, for each observation $i$, the estimated projection vector
@@
We then compute the residuals ${\bf R}_{i}={{\bf Z}}_{i}-\hat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into an $n\times p$ matrix $\mathcal{R}$. Our estimator of the efficient information is $\widehat{I}=\mathcal{R}^{\top}\mathcal{R}/n$, and the asymptotic covariance is $\widehat{I}^{-1}$. In practice this procedure is numerically stable, requires no tuning beyond the main network fit, and produces reliable covariance estimates for $\widetilde{\mbox{${\theta}$}}$.
6 Simulation Study
In this section, we compare the finite-sample performance of the proposed FLEXI-Haz method with the partially linear neural PH estimator of zhong2022deep. Our evaluation focuses on the estimation accuracy of the finite-dimensional parameter $\theta$} and on the coverage properties of the corresponding Wald confidence intervals.
We consider a setting with both nuisance and primary covariates. Let ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where each covariate is independently sampled from a uniform distributed over [1,1][-1,1]. The true parameter of interest is $\mbox{${\theta}$}_{o}=(b_{1},b_{2})^{\top}=(2,-1)^{\top}$. For the nuisance component, we define $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects and interactions among the nuisance covariates. The event time $U$ is generated from a hazard function of the form
$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau]\,.$
@@ 6 Simulation Study
with $\tau=30$ chosen as the fixed study horizon. Right censoring is introduced through an independent censoring time $C$ sampled from an exponential distribution with expectation $1/30$, combined with administration censoring at time 30. The parameters of the data-generating mechanism are chosen so that the censoring rate is approximately 30%.
@@
For both FLEXI-Haz and the estimator of zhong2022deep, the nuisance function is estimated using a fully connected feedforward network with ReLU activations. The only and difference between the two estimators lies in the underlying hazard model, and the training pipeline and tuning strategy are identical. Hyperparameters are selected once, using only one simulated dataset. We construct the expanded dataset using a fixed time grid of $m_{n}=n$ points augmented with all observed times, and perform 5-fold cross-validation over the following search space:
  • number of hidden layers: $\{1,2,3,4,5,6\}$
  • width (nodes per hidden layer): $\{20,40,80,160,320\}$
  • learning rate: $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$
Training uses early stopping with a patience of 35 epochs and a batch size of $10^{5}$ on the expanded dataset. For each hyperparameter combination we train the network once and record the average validation loss across the five folds; the configuration achieving the smallest average loss is selected. Cross-validation selects, for FLEXI-Haz, a network with five hidden layers, width 20, and learning rate $10^{-3}$. For the estimator of zhong2022deep, the selected architecture has four hidden layers, width 80, and the same learning rate $10^{-3}$.
@@
Table 1 presents the results for $n=8000$ based on 200 replications. The proposed estimator, FLEXI-Haz, is essentially unbiased and its empirical variability closely matches the theoretical standard errors, yielding coverage probabilities near their nominal levels. In contrast, the partially linear PH estimator of zhong2022deep exhibits substantial bias in both components of $\boldsymbol{\theta}_{o}$, leading to near-zero empirical coverage.
@@
Table 1: Simulation results for $n=8000$ observations, based on 200 replications. The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$. “Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90% , 95%” the empirical coverage rates of the corresponding Wald-type confidence intervals.
FLEXI-Haz zhong2022deep
Est Emp SD Est SE 90% 95% Est Emp SD Est SE 90% 95%
@@ 7 Discussion
In this work we proposed FLEXI–Haz, a deep semiparametric survival model that combines a fully non-PH formulation with a partially linear structure. By modeling the log-hazard as $\boldsymbol{\theta}_{o}^{\top}\mathbf{Z}+g_{o}(t,\mathbf{X})$ and estimating $g_{o}$ with a deep ReLU network, we retain interpretability and valid inference for the low-dimensional parameter $\boldsymbol{\theta}_{o}$ while allowing highly flexible, time-varying effects of the nuisance covariates. Our main theoretical result is that FLEXI–Haz enjoys the same key properties that zhong2022deep established for partially linear Cox models under PH: the neural estimator of $g_{o}$ attains the optimal composite minimax rate, and the estimator of $\boldsymbol{\theta}_{o}$ is $\sqrt{n}$–consistent, asymptotically normal, and semiparametrically efficient. Crucially, these properties continue to hold in our more general, baseline-free non-PH setting, despite the presence of time in the nuisance component and the use of the full likelihood rather than a partial likelihood.
The simulation study supports these theoretical findings and illustrates their practical relevance. In a setting with nonlinear time–covariate interactions, FLEXI–Haz produces essentially unbiased estimates of $\boldsymbol{\theta}_{o}$, with empirical standard deviations that align closely with the estimated standard errors and nominal coverage for Wald-type confidence intervals. In contrast, the partially linear PH estimator of zhong2022deep, which forces the nuisance component to depend only on $\mathbf{X}$, exhibits substantial bias and severe undercoverage when the true hazard is non-proportional. This behavior is exactly what the theory predicts: when time-varying effects are incorrectly constrained to be time-invariant, the nuisance approximation error leaks into the finite-dimensional parameter. Allowing $g_{o}$ to depend on both $t$ and $\mathbf{X}$ removes this source of misspecification, and the empirical results confirm that the resulting estimator behaves in accordance with the asymptotic theory.
Although FLEXI–Haz provides consistent estimators of the hazard and enables estimation of the conditional survival function $S(t\mid\mathbf{X},\mathbf{Z})$, the present work does not derive the asymptotic distribution of the resulting survival-curve estimator. Nevertheless, valid pointwise confidence intervals and simultaneous confidence bands can still be constructed by applying the ensemble–bootstrap procedure developed by arie2024confidence. That method is fully nonparametric, accommodates right censoring, and can be used as a post–hoc procedure on top of FLEXI–Haz without modifying the training algorithm.
Our implementation focuses on fully connected ReLU networks, which fit naturally into the composite Hölder framework and for which sharp approximation and entropy bounds are available. This choice enables us to obtain clean minimax rates and empirical-process control. At the same time, fully connected architectures are not always the most natural representation for complex structured data such as images. Convolutional neural networks (CNNs) are often preferable in imaging applications, and several deep survival methods in the literature rely on convolutional encoders. While the FLEXI–Haz likelihood and estimation strategy can be used with convolutional backbones in practice, our current theory does not cover convolutional architectures, and the existing results on sparse ReLU networks cannot be directly transferred to CNNs. Extending the approximation and efficiency analysis to convolutional or other structured architectures is therefore an important direction for future work.
A further practical advantage of our formulation is that it meshes naturally with the standard “pseudo-observation” representation used to handle time-dependent covariates. The full-likelihood derivation already requires us to expand each subject into multiple pseudo-observations over risk intervals. This is precisely the data-expansion scheme commonly used to encode time-dependent covariates in Cox-type models. As a result, extending FLEXI–Haz to time-dependent covariates is conceptually straightforward: one augments $\mathbf{X}$ to include time-varying components and constructs the expanded dataset in the usual way. The theoretical development is also compatible with this extension, as the key arguments are expressed in terms of predictable processes and integrals over the at-risk time, and do not rely on time-independence of the covariates beyond notational convenience.
The main current limitation of FLEXI–Haz is computational rather than statistical. Because we work with the full likelihood and approximate the cumulative hazard via a Riemann-type sum on a fine time grid, the expanded dataset contains $O(n\,m_{n})$ pseudo-observations, where $m_{n}$ is the number of grid points. In the default choice $m_{n}\asymp n$, this leads to an effective $O(n^{2})$ cost per epoch, which can become prohibitive for very large studies or dense event processes. Our experiments show that this is manageable for moderate $n$, but improving scalability is crucial for applications with hundreds of thousands of subjects or very fine time resolution.
Several strategies could mitigate this computational burden. One natural approach is to sub-sample pseudo-observations corresponding to non-events in the risk sets, e.g., by keeping all intervals containing failures and a carefully chosen random subset of non-event intervals, with importance weights to preserve unbiasedness. Since the information about $\boldsymbol{\theta}_{o}$ and $g_{o}$ is concentrated near failures and in regions with high hazard, aggressive sub-sampling of “uninformative” background intervals should reduce cost with limited loss in efficiency (Keret02102023). Another line of work is to exploit structured quadrature rules or other powerful approximation methods.
Our theoretical treatment of numerical integration in Theorem 5 is deliberately abstract and suggests a broad class of admissible approximations. The only requirement is that the approximation error in the empirical loss is $O_{p}(n^{-1})$ uniformly over the parameter space, and that the corresponding error in the $\boldsymbol{\theta}$-score is $O_{p}(n^{-1/2})$. Any numerical scheme—whether based on fewer grid points, higher-order quadrature, adaptive time meshes, or stochastic approximations—that satisfies these two error bounds yields the same asymptotic distribution of $\widetilde{\mbox{${\theta}$}}$. In particular, this shows that our choice of a simple Riemann sum with $m_{n}\asymp n$ is sufficient but not necessary: one can potentially use substantially fewer time points while preserving the validity of the asymptotic theory, provided the approximation quality is high enough.
8 Appendix - Proofs

### v2

@@ 6.1 Applying DNN
In practice, we approximate $g_{o}$ by a fully connected ReLU network and model the linear term $\mbox{${\theta}$}^{\top}{\bf Z}$ using a parallel linear layer. The network computes $\widetilde{g}(t,{\bf X})$ and $\widetilde{\mbox{${\theta}$}}^{\top}{\bf Z}$ and their sum is inserted into the loss based on (3.2). This architecture jointly learns the nuisance function $g$ and a finite–dimensional linear effect of $\bf Z$} in a way that mirrors the model structure.
To evaluate the approximated log-likelihood (3.2) in practice, we discretize the time axis on a grid (e.g., $n$) and augment it with all observed times, yielding $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$. The dataset is then expanded in the usual counting‐process manner. Each subject $i$ with data $(T_{i},\Delta_{i},\mathbf{X}_{i},\mathbf{Z}_{i})$ is represented by one dummy-observation for every interval $t_{k},t_{k+1})$ such that $t_{k}\leq T_{i}$. All dummy-observations inherit $(\mathbf{X}_{i},\mathbf{Z}_{i})$; the event indicator $\Delta_{ik}=\mathbf{1}\{T_{i}\in(t_{k-1},t_{k}],\Delta=1\}$ is set to 0 except for the final interval containing $T_{i}$, where it takes the original value $\Delta_{i}$. This choice of grid and data expansion enables efficient computation within standard deep-learning frameworks. The construction parallels the standard expanded-data representation used for Cox models with time-dependent covariates (Therneau and Grambsch, 2000; Hu and Nan, 2023).
@@ 6.2 Estimation of the Covariance Matrix $I(\mbox{${\theta}$}_{o})^{-1}$
Inference on $\mbox{${\theta}$}_{0}$ relies on estimating the efficient information matrix $I(\mbox{${\theta}$}_{o})$, whose inverse provides the asymptotic covariance of the estimator $\widetilde{\mbox{${\theta}$}}$. Recall that the efficient information is
@@
where $\bf g$$^{\ast}$ is the event–weighted projection of $\bf Z$} onto the nuisance tangent space. In practice we estimate $\bf g$$^{\ast}$ using an additional set of neural networks and then construct a plug–in estimate of $I(\mbox{${\theta}$}_{o})$ from the corresponding residuals. Specifically, we approximate $\bf g$$^{\ast}$ coordinatewise. For each $j=1,\ldots,p$ we fit a separate fully connected network $\widehat{g}^{\ast}_{j}(T,{\bf X})$ by least squares using only uncensored observations, that is by $\min\sum_{i:\,\Delta_{i}=1}\{Z_{ij}-\widehat{g}^{\ast}_{j}(T_{i},{{\bf X}}_{i})\}^{2}$. Each network $j$, $j=1,\ldots,p$, takes $(T_{i},\mathbf{X}_{i})$ as input, $Z_{ij}$ as target, and uses the same architecture and hyperparameters as the main model; empirically, these choices produce stable estimates and are not sensitive to tuning.
To avoid overfitting, we employ 55–fold cross–fitting. The data are split into five folds, for each fold and each coordinate $j$, the corresponding network is trained on the remaining four folds (restricting to $\Delta=1$) and predictions $\widehat{g}^{\ast}_{j}(T_{i},X_{i})$ are obtained for the held-out fold. Aggregating across folds yields, for each observation $i$, the estimated projection vector
@@
We then compute the residuals ${\bf R}_{i}=\sqrt{\Delta}_{i}\{{{\bf Z}}_{i}-\widehat{\bf g}^{\ast}(T_{i},{{\bf X}}_{i})\}\in\mathbb{R}^{p}$, $i=1,\ldots,n$, and stack them into a matrix $\mathcal{R}$. Our estimator of the efficient information is $\widehat{I}=\mathcal{R}^{\top}\mathcal{R}/n$, and the asymptotic covariance is $\widehat{I}^{-1}$. In practice this procedure is numerically stable, requires no tuning beyond the main network fit, and produces reliable covariance estimates for $\widetilde{\mbox{${\theta}$}}$.
7 Simulation Study and Real-data application
7.1 Simulation Study
We compare the finite-sample performance of the proposed FLEXI–Haz method with the partially linear neural PH estimator of Zhong et al. (2022). Our focus is on estimation of the finite-dimensional parameter $\theta$} and on the coverage of the associated Wald confidence intervals. We consider nuisance and primary covariates ${{\bf X}}=(X_{1},X_{2},X_{3})^{\top}$ and ${{\bf Z}}=(Z_{1},Z_{2})^{\top}$, where all covariates are independently distributed as $\mathrm{Unif}[-1,1]$. The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$. The nuisance component includes $f({\bf X})=0.2(X_{1}+X_{2})+0.5X_{1}X_{2}+X_{3}^{2}$, which induces nonlinear and interaction effects. Event times are generated from
$h_{o}(t\mid{\bf X},{\bf Z})=0.1\exp\!\left\{(0.1+f({\bf X})^{2})t+\mbox{${\theta}$}_{o}^{\top}{\bf Z}\right\},\qquad t\in[0,\tau],$
@@ 7.1 Simulation Study
with fixed study horizon $\tau=30$. Right censoring is introduced through an independent exponential censoring time with mean $1/30$, together with administrative censoring at time 30. The resulting censoring rate is approximately 30%.
@@
For both FLEXI–Haz and the model of Zhong et al. (2022), the nuisance function is estimated using a fully connected feedforward ReLU network. The two methods differ only in the underlying hazard model; the training pipeline and tuning strategy are otherwise identical. Hyperparameters are selected once using a single simulated dataset. We construct the expanded dataset on a grid of $m_{n}=n$ time points augmented with all observed times, and perform 5-fold cross-validation over the number of hidden layers $\{1,2,3,4,5,6\}$, width $\{20,40,80,160,320\}$, and learning rate $\{10^{-3},\,5\times 10^{-3},\,10^{-2},\,5\times 10^{-2}\}$. Training uses early stopping with patience 35 and batch size $10^{5}$ on the expanded dataset. For each hyperparameter combination, the network is trained once and the average validation loss across the five folds is recorded; the configuration with the smallest average loss is selected. Cross-validation selects, for FLEXI–Haz, five hidden layers, width 20, and learning rate $10^{-3}$, and for Zhong et al. (2022), four hidden layers, width 80, and learning rate $10^{-3}$.
@@
Table 1 reports results for $n=8000$ based on 200 replications. FLEXI–Haz is essentially unbiased, and its empirical standard deviation closely matches the average estimated standard error, yielding coverage probabilities close to the nominal levels. In contrast, the partially linear PH estimator of Zhong et al. (2022) exhibits substantial bias in both components of $\mbox{${\theta}$}_{o}$, resulting in near-zero empirical coverage.
To assess the finite-sample performance of the proposed pointwise survival-curve inference, we conducted an additional simulation experiment under the same setup as above, except that each nuisance covariate was generated independently from a Poisson(2)(2) distribution truncated at $6$. In each of $200$ simulation runs, we generated a training sample of size $n=8000$ and an additional evaluation sample of size $200$. For each target level $s\in\{0.1,0.2,\ldots,0.9\}$, we selected a fixed time point $t_{s}$ such that the mean estimated survival in the training sample was approximately $s$, and evaluated the empirical coverage of the pointwise confidence interval for the survival probability at time $t_{s}$ based on the evaluation samples.
Table 2 demonstrates that the proposed pointwise inference procedure is well calibrated in finite samples. Across the selected time points, the empirical coverage remains close to the nominal level for both the 90% and 95% confidence intervals, supporting the practical reliability of the method in the discrete-covariate setting considered here. The slight overcoverage at $t=6.7126$ may be due to boundary effects, since this is a late time point where the true survival probability is close to zero for some values of $({\bf X}_{0},{\bf Z}_{0})$.
@@
Table 1: Simulation results for $n=8000$ observations, based on 200 replications. The true parameter is $\mbox{${\theta}$}_{o}=(2,-1)^{\top}$. “Est” denotes the mean of the estimated coefficients; “Emp SD” the empirical standard deviation; “Est SE” the estimated theoretical standard error; and “90%CI , 95%CI” the empirical coverage rates of the corresponding Wald-type confidence intervals.
FLEXI–Haz Zhong et al. (2022)
Est Emp SD Est SE 90%CI 95%CI Est Emp SD Est SE 90%CI 95%CI
@@
Table 2: Simulation results of empirical pointwise coverage of the 90% and 95% confidence intervals for the survival curve, based on 200200 replications with training sample size $n=8000$ and testing sample size of $200$. For each target survival time $t_{s}$, we evaluated whether the confidence interval covered the true survival value at $t_{s}$.
Marginal survival $s$ Target $t_{s}$ 90%CI SD of coverage 95%CI SD of coverage
0.1 6.7126 0.9377 0.0741 0.9676 0.0531
0.2 4.2934 0.9120 0.0861 0.9548 0.0638
0.3 3.0192 0.9167 0.0769 0.9565 0.0595
0.4 2.2003 0.9121 0.0755 0.9595 0.0523
0.5 1.6035 0.8903 0.0889 0.9488 0.0589
0.6 1.2330 0.8997 0.0912 0.9520 0.0611
0.7 0.9476 0.8995 0.0824 0.9480 0.0614
0.8 0.6812 0.9123 0.0815 0.9577 0.0559
0.9 0.3900 0.9044 0.0987 0.9567 0.0638
7.2 Real-data application: Rotterdam breast cancer study
To illustrate the practical performance of FLEXI–Haz, we analyzed the Rotterdam Breast Cancer data set, available in the survival package in R. These data were studied by Foekens et al. (2000) to assess clinical factors associated with time to breast cancer recurrence or death, where the outcome is defined as the number of days from primary surgery until the earlier of recurrence or death. The data set contains $n=2982$ patients, with an overall censoring proportion of $57.35\%$, and includes nine baseline covariates: age, progesterone receptor level, estrogen receptor level, number of positive lymph nodes, menopausal status, tumor size, tumor grade, chemotherapy, and hormonal therapy. Among these variables, the first four are continuous and the remaining five are discrete.
Our primary scientific goal in this analysis is to assess the effects of chemotherapy and hormonal therapy. Accordingly, we included these two treatment indicators in the linear component $\bf Z$}, while modeling the remaining covariates nonparametrically through the DNN component.
We used five-fold cross-validation. Within each fold, the training portion was further split so that $20\%$ of the training data served as a validation set for tuning hyperparameters, after which model performance was evaluated on the held-out test set. Table 3 summarizes the estimated treatment coefficients and corresponding $95\%$ confidence intervals for the proposed method, together with the comparator methods.
Table 3: Estimated treatment effects and $95\%$ confidence intervals for the Rotterdam breast cancer data.
Treatment Method Estimate (SE) 95% CI
Chemotherapy Cox PH 0.127(0.071)-0.127\ (0.071) (0.267, 0.013)(-0.267,\ 0.013)
Zhong et al. (2022) 0.257(0.095)-0.257\ (0.095) (0.443,0.071)(-0.443,\ -0.071)
FLEXI–Haz 0.296(0.094)-0.296\ (0.094) (0.480,0.112)(-0.480,\ -0.112)
Hormonal therapy Cox PH 0.103(0.083)-0.103\ (0.083) (0.266, 0.061)(-0.266,\ 0.061)
Zhong et al. (2022) 0.372(0.080)-0.372\ (0.080) (0.529,0.215)(-0.529,\ -0.215)
FLEXI–Haz 0.365(0.080)-0.365\ (0.080) (0.522,0.209)(-0.522,\ -0.209)
Table 3 shows negative estimated effects of both chemotherapy and hormonal therapy, with 95% confidence intervals excluding zero under FLEXI–Haz. The corresponding estimates are qualitatively similar to those of Zhong et al. (2022), whereas the Cox PH model yields confidence intervals that include zero for both treatments. These findings indicate that FLEXI–Haz recovers interpretable treatment effects while accommodating flexible nonproportional nuisance effects.
Figure 1 displays estimated survival curves for 16 randomly selected subjects. The plug-in estimator is visually smooth because it is evaluated on a dense time grid, whereas the one-step estimator is piecewise constant, reflecting that its bias-correction term is constructed from subjects sharing the same discrete nuisance-covariate profile ${\bf X}_{0}$.
Figure 1: Estimated survival curves for 16 randomly selected subjects in the Rotterdam breast cancer data. The blue curve is the plug-in FLEXI–Haz estimator, evaluated on a dense grid. The red step function is the cross-fitted one-step estimator, and the shaded band gives the corresponding pointwise confidence interval. The visible steps in the one-step curves arise from the finite-support correction, which uses subjects with the same nuisance-covariate profile ${\bf X}_{0}$.
@@ 8 Discussion
We proposed FLEXI–Haz, a deep semiparametric survival model that combines a partially linear structure with a flexible non-proportional hazards formulation. By modeling the log-hazard as $\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(t,{\bf X})$, the method preserves interpretability for the low-dimensional parameter $\mbox{${\theta}$}_{o}$ while allowing the nuisance component to vary nonlinearly with both time and covariates. Our main theoretical results show that the DNN estimator of $g_{o}$ attains the optimal composite minimax rate, that the estimator of $\mbox{${\theta}$}_{o}$ is $\sqrt{n}$-consistent, asymptotically normal, and semiparametrically efficient, and that pointwise frequentist asymptotic inference is available for the survival curve of a new subject through a cross-fitted one-step procedure. The simulation and real-data analyses support these conclusions and illustrate the practical advantages of FLEXI–Haz when proportional hazards may be too restrictive.
The main limitation of the current survival-inference theory is the additional assumption that the nuisance covariate vector $\bf X$} has discrete finite support, which is used to obtain an explicit Riesz representer and a tractable one-step correction. Although this restriction is often reasonable in applications with categorical covariates or grouped continuous features, extending the pointwise survival-inference theory to more general covariate distributions is an important direction for future work. Another natural extension is to develop analogous theory for richer network architectures beyond fully connected ReLU networks.
Appendix A Proofs for Section 4